OpenStudy (anonymous):

Multiply (5-3i)(6+4i) Simplify i^77 radical -60

OpenStudy (anonymous):

The first one, multiply the binomial the way you are used to do it, and in the end substitute for \[ i^2 = -1\]

mathslover (mathslover):

Instead of solving this question lemme ask you one thing : what is (x-1)(y+2) ?

OpenStudy (anonymous):

second one is just playing with cyclic expressions \[ i^2 = -1 \\i^3=-i \\i^4=1\]

mathslover (mathslover):

@kenya.porter .. ?

OpenStudy (anonymous):

I wont (-: Fortunately there are no solutions mentioned yet, only ways.

mathslover (mathslover):

:) @kenya.porter please reply ..

OpenStudy (anonymous):

xy+2x-y-2

mathslover (mathslover):

very good .. so you did like this : \[\large{(x-1)(y+2)=x(y)+x(2)-1(y)-1(2)=xy+2x-y-2}\] right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok so apply the same process in the given question .. \[\large{(5-3i)(6+4i)=?}\] can u tell me what r u getting?

OpenStudy (anonymous):

42+2i

mathslover (mathslover):

\[\large{30+20i-18i-12i^2}\] \[\large{30+2i+12}\] \[\large{42+2i}\] very good so this is your answer ..

mathslover (mathslover):

so shall i move on to second one @kenya.porter ??

OpenStudy (anonymous):

yes,thank you so much i would appreciate it.

mathslover (mathslover):

\[\large{i^{77}}\] \[\large{(i^7)^{11}}\] \[\large{i^7=i^3*i^4}\] \[\large{i^7=-i * 1}\] \[\large{i^7=-i}\] hence \[\large{(i^7)^{11}=(-i)^{11}}\] \[\large{(-i)^{11}=(-1)*(i)^{11}=-1*i^3*i^4*i^4}\] \[\large{-1*-i*1*1}\] \[\large{i}\]

mathslover (mathslover):

got it @kenya.porter ?

OpenStudy (anonymous):

yes i do

mathslover (mathslover):

any more problems? @kenya.porter

OpenStudy (anonymous):

yes radical -60

mathslover (mathslover):

ok !! : \[\large{i^{-60}}\] \[\large{\frac{1}{i^{60}}}\] \[\large{\frac{1}{(i^5*i^{10})^4}}\] \[\large{\frac{1}{(i*-i)^4}}\] \[\large{\frac{1}{(-i)^4}}\] \[\large{\frac{1}{1}}\] \[\large{ 1}\]

mathslover (mathslover):

Introduction to complex numbers : http://openstudy.com/study#/updates/4fe196d3e4b06e92b870a14a Practicing on complex numbers : http://openstudy.com/study#/updates/4fe5a252e4b06e92b873c296 best of luck..

OpenStudy (anonymous):

ok thank you